# Not All Rollouts are Useful: Down-Sampling Rollouts in LLM Reinforcement Learning

### v3

@@ Abstract
Reinforcement learning with verifiable rewards (RLVR) has emerged as the leading approach for enhancing reasoning capabilities in large language models. However, it faces a fundamental compute and memory asymmetry: rollout generation is embarrassingly parallel and memory-light, whereas policy updates are communication-heavy and memory-intensive. To address this, we introduce PODS (Policy Optimization with Down-Sampling), which decouples rollout generation from policy updates by training only on a strategically selected subset of rollouts, maintaining learning quality while dramatically reducing update costs. We propose a principled subset selection criterion—max-variance down-sampling—that maximizes reward diversity, and provide an efficient $O(n\log n)$ implementation. Empirically, Group Relative Policy Optimization (GRPO) with PODS achieves the peak test accuracy of vanilla GRPO at least $\mathbf{1.7\times}$ faster across the different reasoning benchmarks and hardware configurations we tested.
@@ 1 Introduction
We address this bottleneck through a key observation: not all rollouts contribute equally to model improvement. Beyond a certain scale, additional rollouts provide diminishing returns and can even degrade learning signals through redundant information. This suggests a natural solution: generate large batches of rollouts during the scalable inference phase, but train selectively on only the most informative subset during the policy update phase, avoiding the latency overhead of memory-saving techniques. We formalize this idea in PODS (Policy Optimization with Down-Sampling). As illustrated in Fig.˜2, PODS maximizes hardware utilization by generating $n$ rollouts per prompt but updating on only $m<n$ informative samples selected by a principled down-sampling rule.
Within the PODS framework, we introduce max-variance down-sampling, a principled criterion that selects the subset of rollouts with the greatest reward variance of the selected subset, thereby preserving strong contrastive signals. We show that the resulting combinatorial problem can be solved in $O(n\log n)$ time and, in the common binary-reward setting, reduces to picking the $m/2$ highest-reward and $m/2$ lowest-reward rollouts. We evaluate PODS with GRPO on GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021) across multiple model and hardware configurations, demonstrating that it achieves the peak test accuracy of baseline GRPO at least $\mathbf{1.7\times}$ faster.
Figure 1: Inference scales efficiently while policy updates become memory-bound in RLVR. Empirical timing breakdown when fine-tuning Qwen2.5-3B-Instruct on GSM8K using 88 A100-80GB GPUs with varying rollouts per GPU. Top: Total wall-clock time per iteration. Policy updates hit memory limits after $32$ rollouts per GPU (OOM beyond this point), requiring gradient accumulation that dramatically slows training. Bottom: Per-token inference time decreases $21\times$ through batching (from $8$ to $512$ rollouts), saturating beyond $512$. This demonstrates the core asymmetry that PODS exploits: inference parallelizes efficiently while policy updates become memory-bound.
Figure 2: Visualization of three training strategies: vanilla GRPO, GRPO with gradient accumulation (GRPO-GA), and GRPO with PODS (GRPO-PODS). Vanilla GRPO generates $n$ rollouts and trains on all of them, leaving inference hardware underutilized. GRPO-GA alleviates this issue with memory-saving techniques such as gradient accumulation, but at the cost of more sequential steps in the policy-update phase. In contrast, GRPO-PODS also generates $n$ rollouts but trains on only $m$ carefully selected examples, maximizing inference utilization, avoiding gradient-accumulation overhead, and providing a cleaner learning signal that yields better final performance.
@@ 2 Related Work
Reinforcement learning has emerged as a powerful paradigm for enhancing the reasoning capabilities of LLMs across math, coding, and problem-solving domains (Jaech et al., 2024; Shao et al., 2024; Kazemnejad et al., 2024). Although classical algorithms such as Proximal Policy Optimization (PPO) (Schulman et al., 2017) laid the foundation, recent work has tailored them specifically to language models. Group Relative Policy Optimization (GRPO) (Shao et al., 2024) has gained prominence for reasoning tasks because of its implementation simplicity, competitive performance relative to PPO, and lack of a separate critic network. OpenAI o1 (OpenAI, 2024) and DeepSeek R1 (Guo et al., 2025), which used large-scale RL, have sparked interest in reasoning-focused RL methods (Chen et al., 2025; Hu et al., 2025; Hu, 2025; Cui et al., 2025). Meanwhile, value-based approaches like PPO remain central (Yuan et al., 2025a; b), alongside complementary techniques such as Monte Carlo Tree Search (Gao et al., 2024; Xie et al., 2024) and multi-agent methods (FAIR et al., 2022). A recent line of work has also explored data selection for improving RL methods for LLM training. Specifically, prompt selection and filtering has gained significant attention from works such as DAPO (Yu et al., 2025), SRPO (Zhang et al., 2025), Reinforce-Rej (Xiong et al., 2025), Polaris An et al. (2025) and VAPO (Yuan et al., 2025a). Our method advances this line of work by focusing on down-sampling rollouts within each prompt, instead of selecting or filtering prompts themselves. By tackling this computational-efficiency bottleneck, our approach complements existing methods and can be combined with them to further improve reasoning performance.
@@
The scale of modern machine learning necessitates effective data management strategies, particularly as datasets grow larger, noisier, and more imbalanced. Training on the full dataset can be prohibitively expensive, motivating sophisticated data-selection and down-sampling methods. Such techniques succeed across diverse settings—from theoretical results in clustering (Har-Peled & Mazumdar, 2004), regression (Li et al., 2013; Rudelson & Vershynin, 2007; Clarkson, 2010) to practical systems in speech recognition (Liu et al., 2015; Wei et al., 2014) and computer vision (Kaushal et al., 2019; Bankes et al., 2024). In reinforcement learning, prioritized experience replay (Schaul et al., 2015) and related methods (Hou et al., 2017; Saglam et al., 2023; Cusumano-Towner et al., 2025) highlight the value of selective sampling from experience buffers. More recently, careful data selection has become central to foundation-model training (Goyal et al., 2024; Schuhmann et al., 2021; Gadre et al., 2023) and emerging applications such as computational advertising (Bei et al., 2023; Gravin et al., 2024). Yet, to our knowledge, we are the first to apply principled down-sampling to the rollout-generation stage of LLM reinforcement learning, mitigating a key computational bottleneck while strengthening the learning signal.
@@ 3 Down-Sampling Rollouts in GRPO
In this section, we present our approach to resolving the computational asymmetry between inference and policy updates in LLM reinforcement learning. We first review the original GRPO algorithm in Section˜3.1, highlighting its structural components and computational demands. Next, in Section˜3.2, we introduce the PODS (Policy Optimization with Down-Sampling) framework, which strategically selects informative rollouts to maximize hardware utilization during both inference and policy-update phases. In Section˜3.3, we develop a principled max-variance down-sampling method that preserves strong contrastive signals, justified by Razin et al. (2025), by retaining only rollouts from the extremes of the reward spectrum. We show that this method admits an elegant, $O(n\log n)$ solution, making it practical for real-world deployment. Overall, our framework retains the advantages of GRPO while boosting computational and memory efficiency across diverse hardware setups.
@@ 3.1 Preliminaries
Let $\pi_{\theta}$ denote the policy parameterized by $\theta$, which defines a distribution over next-token probabilities given the previous tokens in a sequence. Given a single input prompt $p$ (e.g., a math problem), GRPO first generates a group of $n$ rollouts $\mathbf{o}=(o_{1},o_{2},\dots,o_{n})$} by autoregressively sampling from $\pi_{\theta}$. Each rollout is a complete token sequence excluding the prompt, representing a possible solution. Each rollout is then evaluated using a reward model $r_{i}=R(o_{i})$, which scores the quality and correctness of the corresponding output $o_{i}$. This yields a reward vector $\mathbf{r}=(r_{1},r_{2},\dots,r_{n})$. From these rewards, we compute normalized advantage estimates: $a_{i}={(r_{i}-\mu)}/{\sigma}$, where $\mu$ and $\sigma$ are the mean and standard deviation of the rewards respectively.
@@ 3.2 PODS Framework
$D(\mathbf{o},\mathbf{r};m)$ is a function that takes as inputs $n$ rollouts $\mathbf{o}=(o_{1},o_{2},\ldots,o_{n})$, their corresponding rewards $\mathbf{r}=(r_{1},r_{2},\ldots,r_{n})$, and the update size $m$. It outputs a subset of indices $S\subseteq\{1,2,\dots,n\}$, where $|S|=m$, indicating which rollouts to retain for the policy update phase.
@@ 3.3 Max-Variance Down-Sampling
Let $\mu$ be the mean of the rewards in $S^{*}$. Since the set $S^{*}$ does not take the form of $\{1,\dots,{m-k}\}\cup\{{n-k+1},\dots,n\}$} for any $k$, there exists either (i) an element $i\in S^{*}$ such that $i>1,r_{i}\leq\mu$ and $i-1\not\in S^{*}$, or (ii) an element $j\in S^{*}$ such that $j<n,r_{j}\geq\mu$ and $j+1\not\in S^{*}$. That is, there exists an element in $S^{*}$, such that another element further from $\mu$ is not in $S^{*}$. We will show that we can swap them without decreasing variance.
@@
Lemma˜3.1 naturally leads to a practical algorithm, Algorithm˜2, for max-variance down-sampling. Moreover, it also offers intuition as to why maximizing variance is effective: the optimal subset contains the $k$ highest rewards and the (mk)(m-k) lowest rewards, thereby capturing strong contrastive signals from both positive and negative examples.
@@ 4 Experiments
We evaluate PODS across diverse hardware configurations, model architectures, and model scales to demonstrate its generalizability and practical benefits. We test on two mathematical reasoning benchmarks—GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021)—using Qwen2.5 (Qwen et al., 2025) and Llama3.2 (MetaAI, 2024) models ranging from 3B to 7B parameters. Our experimental design covers both resource-constrained single-GPU setups and multi-GPU distributed training to validate PODS across different deployment scenarios. Table˜1 describes our experimental configurations. To facilitate reproduction of our results, we publish the code for our experiments at https://github.com/YixuanEvenXu/pods.
@@
(b) MATH Qwen2.5 3B 1 L40S LoRA (rank 64, $\alpha=64$)
(c) GSM8K Llama3.2 3B 1 L40S LoRA (rank 64, $\alpha=64$)
(d) GSM8K Qwen2.5 3B 8 H100s Full-Parameter
(e) GSM8K Qwen2.5 7B 8 A100s Full-Parameter
Table 1: Experimental configurations testing PODS across model scales, hardware constraints, and training paradigms. Settings (a-c) test resource-constrained scenarios with LoRA fine-tuning, while (d-e) evaluate full-parameter training with distributed setups.
@@
For settings (a-c), we use Unsloth (Daniel Han & team, 2023) with TRL (von Werra et al., 2020) for efficient LoRA (Hu et al., 2022) fine-tuning. For settings (d-e), we implement distributed training with DeepSpeed ZeRO-2 (Rajbhandari et al., 2020) and extend the open-r1 library (HuggingFace, 2025) to support PODS on multiple devices.
@@
We employ rule-based reward models that score rollouts, following standard practices in mathematical reasoning evaluation. Specifically, we reward an answer for correctness, format compliance, and the right number of thinking tags separately, resulting in a discrete but non-binary reward function. Details are provided in Section˜A.1.
@@
In Section˜4.1, we compare the performance of GRPO and GRPO-PODS across five hardware and model settings listed in Table˜1. We show that for all the settings we test, GRPO-PODS consistently outperforms GRPO in terms of performance as the training time increases. Then, in Section˜4.2, we focus on setting (a), and analyze the effect of the rollout and update sizes $(n,m)$ on the performance of GRPO-PODS, providing empirical insights into how to choose the rollout and update sizes for GRPO-PODS. We present additional experiments about different down-sampling rules, and evaluation results about PODS’s speed up ratio compared to GRPO and the average response length over the course of training in Sections˜A.3, A.4 and A.5, respectively.
@@ 4.1 Comparing GRPO-PODS to baseline GRPO
We evaluate max-variance down-sampling PODS with GRPO against baseline GRPO using two experimental designs reflecting real-world constraints. For single-GPU settings (a-c), we compare against vanilla GRPO with matched training batch sizes $(m)$, where $m$ is selected to fit within memory. This corresponds to the comparison between GRPO and GRPO-PODS in Fig.˜2. For distributed settings (d-e), we compare against GRPO with gradient accumulation (GRPO-GA)—the standard approach for scaling RLVR. In GRPO-GA, large batches are processed through multiple gradient accumulation steps, enabling updating on larger effective batch sizes at the cost of increased communication overhead and iteration time. We fix the total rollouts generated per prompt $(n)$ and compare GRPO-GA against GRPO-PODS. The detailed hyperparameters used for our experiments are listed in Section˜A.2. Results across all five configurations show consistent improvements.
Fig.˜3 shows test accuracy over wall-clock training time across all configurations. PODS consistently achieves faster convergence: reaching the baselines’ peak accuracies at least $1.7\times$ faster (see Section˜A.4 for complete results) while often exceeding final performance. These results demonstrate PODS’s broad applicability across model scales (3B-7B), architectures (Qwen2.5, Llama3.2), and deployment scenarios, making it a practical improvement for RLVR systems using GRPO.
(a) Training Qwen2.5 (3B) on GSM8K with one L40S GPU
(b) Training Qwen2.5 (3B) on MATH with one L40S GPU
(c) Training Llama3.2 (3B) on GSM8K with one L40S GPU
(d) Training Qwen2.5 (3B) on GSM8K with 8 H100 GPUs
(e) Training Qwen2.5 (7B) on GSM8K with 8 A100 GPUs

### v4

@@ Abstract
Reinforcement learning with verifiable rewards (RLVR) has emerged as the leading approach for enhancing reasoning capabilities in large language models. However, it faces a fundamental compute and memory asymmetry: rollout generation is embarrassingly parallel and memory-light, whereas policy updates are communication-heavy and memory-intensive. To address this, we introduce PODS (Policy Optimization with Down-Sampling), which decouples rollout generation from policy updates by training only on a strategically selected subset of rollouts, maintaining learning quality while dramatically reducing update costs. We propose a principled subset selection criterion—max-variance down-sampling—that maximizes the variance of reward in the selected subset, and provide an efficient $O(n\log n)$ implementation of this rule. Empirically, Group Relative Policy Optimization (GRPO) coupled with PODS achieves the peak test accuracy of vanilla GRPO at least $\mathbf{1.7\times}$ faster across the different reasoning benchmarks and hardware configurations we tested.
@@ 1 Introduction
A key observation of our work is: not all rollouts contribute equally to model improvement. Beyond a certain scale, additional rollouts provide diminishing returns and can even degrade learning signals through redundant information. This suggests a natural solution: generate large batches of rollouts during the scalable inference phase, but train selectively on only the most informative subset during the policy update phase, avoiding the latency overhead of memory-saving techniques. We formalize this idea in PODS (Policy Optimization with Down-Sampling). As illustrated in Fig.˜2, PODS maximizes hardware utilization by generating $n$ rollouts per prompt but updating on only $m<n$ informative samples selected by a principled down-sampling rule.
Within the PODS framework, we introduce max-variance down-sampling, a principled criterion that selects the subset of rollouts with the greatest reward variance of the selected subset, thereby preserving strong contrastive signals. We show that the resulting combinatorial problem can be solved in $O(n\log n)$ time and, in the common binary-reward setting, reduces to picking the $m/2$ highest-reward and $m/2$ lowest-reward rollouts. We evaluate PODS with GRPO on GSM8K (Cobbe et al., 2021), MATH (Hendrycks et al., 2021) and the Chemistry subset of SciKnowEval (Feng et al., 2024) across multiple model and hardware configurations, demonstrating that it achieves the peak test accuracy of baseline GRPO at least $\mathbf{1.7\times}$ faster.
Figure 1: Inference scales efficiently while policy updates become memory-bound in RLVR. Empirical timing breakdown when fine-tuning Qwen2.5-3B-Instruct on GSM8K using 88 A100-80GB GPUs with varying rollouts per GPU. Top: Total wall-clock time per iteration. Policy updates hit memory limits after $32$ rollouts per GPU (Out of memory beyond this point), requiring gradient accumulation that dramatically slows training. Bottom: Per-token inference time decreases $21\times$ through batching (from $8$ to $512$ rollouts), saturating beyond $512$. This demonstrates the core asymmetry that PODS exploits: inference parallelizes efficiently while policy updates become memory-bound.
Figure 2: Visualization of three training strategies: vanilla GRPO, GRPO with gradient accumulation (GRPO-GA), and GRPO with PODS (GRPO-PODS). Vanilla GRPO generates $n$ rollouts and trains on all of them, leaving inference hardware underutilized due to the asymmetric computational demands of the two phases. GRPO-GA alleviates this issue with memory-saving techniques such as gradient accumulation, but at the cost of more sequential steps in the policy-update phase. In contrast, GRPO-PODS also generates $n$ rollouts but trains on only $m$ carefully selected examples, maximizing inference utilization, avoiding gradient-accumulation overhead, and providing a cleaner learning signal that yields better final performance.
@@ 2 Related Work
Reinforcement learning has emerged as a powerful paradigm for enhancing the reasoning capabilities of LLMs across math, coding, and problem-solving domains (Jaech et al., 2024; Shao et al., 2024; Kazemnejad et al., 2024). Although classical algorithms such as Proximal Policy Optimization (PPO) (Schulman et al., 2017) laid the foundation, recent work has tailored them specifically to language models. Group Relative Policy Optimization (GRPO) (Shao et al., 2024) has gained prominence for reasoning tasks because of its implementation simplicity, competitive performance relative to PPO, and lack of a separate critic network. OpenAI o1 (OpenAI, 2024) and DeepSeek R1 (Guo et al., 2025), which used large-scale RL, have sparked interest in reasoning-focused RL methods (Chen et al., 2025; Hu et al., 2025; Hu, 2025; Cui et al., 2025). Meanwhile, value-based approaches like PPO remain central (Yuan et al., 2025a; b), alongside complementary techniques such as Monte Carlo Tree Search (Gao et al., 2024; Xie et al., 2024) and multi-agent methods (FAIR et al., 2022). A recent line of work has also explored data selection for improving RL methods for LLM training. Specifically, prompt selection and filtering has gained significant attention from works such as DAPO (Yu et al., 2025), SRPO (Zhang et al., 2025), Reinforce-Rej (Xiong et al., 2025), Polaris (An et al., 2025), GRESO (Zheng et al., 2025) and VAPO (Yuan et al., 2025a). Our method advances this line of work by focusing on down-sampling rollouts within each prompt, instead of selecting or filtering prompts themselves. By tackling this computational-efficiency bottleneck, our approach complements existing methods and can be combined with them to further improve reasoning performance.
@@
The scale of modern machine learning necessitates effective data management strategies, particularly as datasets grow larger, noisier, and more imbalanced. Training on the full dataset can be prohibitively expensive, motivating sophisticated data-selection and down-sampling methods. Such techniques succeed across diverse settings—from theoretical results in clustering (Har-Peled and Mazumdar, 2004), regression (Li et al., 2013; Rudelson and Vershynin, 2007; Clarkson, 2010) to practical systems in speech recognition (Liu et al., 2015; Wei et al., 2014) and computer vision (Kaushal et al., 2019; Bankes et al., 2024). In reinforcement learning, prioritized experience replay (Schaul et al., 2015) and related methods (Hou et al., 2017; Saglam et al., 2023; Cusumano-Towner et al., 2025) highlight the value of selective sampling from experience buffers. More recently, careful data selection has become central to foundation-model training (Goyal et al., 2024; Schuhmann et al., 2021; Gadre et al., 2023) and emerging applications such as computational advertising (Bei et al., 2023; Gravin et al., 2024). Yet, to our knowledge, we are the first to apply principled down-sampling to the rollout-generation stage of LLM reinforcement learning, mitigating a key computational bottleneck while strengthening the learning signal.
@@ 3 Down-Sampling Rollouts in GRPO
In this section, we present our approach to resolving the computational asymmetry between inference and policy updates in LLM reinforcement learning. We first review the original GRPO algorithm in Section˜3.1, highlighting its structural components and computational demands. Next, in Section˜3.2, we introduce the PODS (Policy Optimization with Down-Sampling) framework, which strategically selects informative rollouts to maximize hardware utilization during both inference and policy-update phases. In Section˜3.3, we develop a principled max-variance down-sampling method that preserves strong contrastive signals, justified by Razin et al. (2025), by retaining only rollouts from the extremes of the reward spectrum. We show that this method admits an elegant, $O(n\log n)$ solution, making it practical for real-world deployment.
@@ 3.1 Preliminaries
Let $\pi_{\theta}$ denote the policy parameterized by $\theta$, which defines a distribution over next-token probabilities given the previous tokens in a sequence. Given a single input prompt $p$ (e.g., a math problem), GRPO first generates a group of $n$ rollouts $\mathbf{o}=(o_{1},o_{2},\dots,o_{n})$} by autoregressively sampling from $\pi_{\theta}$. Each rollout is a complete token sequence excluding the prompt, representing a possible solution. Each rollout is then evaluated using a reward model $r_{i}=R(o_{i})$, which scores the quality and correctness of the corresponding output $o_{i}$. This yields a reward vector $\mathbf{r}=(r_{1},r_{2},\dots,r_{n})$. We then compute normalized advantage estimates: $a_{i}={(r_{i}-\mu)}/{\sigma}$, where $\mu$ and $\sigma$ are the mean and standard deviation of the rewards respectively.
@@ 3.2 PODS Framework
$D(\mathbf{o},\mathbf{r};m)$ is a function that takes $n$ rollouts $\mathbf{o}=~(o_{1},o_{2},\ldots,o_{n})$, their corresponding rewards $\mathbf{r}=(r_{1},r_{2},\ldots,r_{n})$, and the update size $m$. It outputs a subset of indices $S\subseteq\{1,2,\dots,n\}$, where $|S|=m$, indicating which rollouts to retain for the policy update phase.
@@
Percentile down-sampling.
The rule $D_{\mathrm{perc}}$ selects the $m$ rollouts by choosing the $\frac{0.5}{m}$, $\frac{1.5}{m}$, $\dots$, $\frac{m-0.5}{m}$ quantiles of the reward distribution. This method ensures that the selected rollouts are evenly spaced across the reward spectrum, providing a more representative sample than random down-sampling.
@@ 3.3 Max-Variance Down-Sampling
Let $\mu$ be the mean of the rewards in $S^{*}$. Since $S^{*}$ does not take the form of $\{1,\dots,{m-k}\}\cup\{{n-k+1},\dots,n\}$} for any $k$, there exists either (i) an element $i\in S^{*}$ such that $i>1,r_{i}\leq\mu$ and $i-1\not\in S^{*}$, or (ii) an element $j\in S^{*}$ such that $j<n,r_{j}\geq\mu$ and $j+1\not\in S^{*}$. That is, there exists an element in $S^{*}$, such that another element further from $\mu$ is not in $S^{*}$. We will show that we can swap them without decreasing variance.
@@
Lemma˜3.1 naturally leads to a practical algorithm, Algorithm˜2, for max-variance down-sampling. Moreover, it also offers intuition as to why maximizing variance is effective: the optimal subset contains the $k$ highest and the (mk)(m-k) lowest rewards, capturing contrastive signals from both positive and negative examples.
@@ 4 Experiments
We evaluate PODS across diverse hardware configurations, model architectures, and model scales to demonstrate its generalizability and practical benefits. We test on three reasoning benchmarks of different domains, GSM8K (Cobbe et al., 2021), MATH (Hendrycks et al., 2021) and the L3 Chemistry subset of SciKnowEval (Feng et al., 2024), using Qwen2.5 (Qwen et al., 2025) and Llama3.2 (MetaAI, 2024) models ranging from 3B to 7B parameters. Our experimental design covers both resource-constrained single-GPU setups and multi-GPU distributed training to validate PODS across different deployment scenarios. Table˜1 describes our experimental configurations. We publish the code for our experiments at https://github.com/YixuanEvenXu/pods.
@@
(b) GSM8K Llama3.2 3B 1 L40S LoRA (rank 64, $\alpha=64$)
(c) MATH Qwen2.5 3B 1 L40S LoRA (rank 64, $\alpha=64$)
(d) Chemistry Qwen2.5 3B 1 L40S LoRA (rank 64, $\alpha=64$)
(e) GSM8K Qwen2.5 3B 8 H100s Full-Parameter
(f) GSM8K Qwen2.5 7B 8 A100s Full-Parameter
Table 1: Experimental configurations testing GRPO with PODS across task domains, model scales, hardware constraints, and training paradigms. Settings (a-d) test resource-constrained scenarios with LoRA fine-tuning, while (e-f) evaluate full-parameter training with distributed setups.
@@
For settings (a-d), we use Unsloth (Daniel Han and team, 2023) with TRL (von Werra et al., 2020) for efficient LoRA (Hu et al., 2022) reinforcement learning. For settings (e-f), we implement distributed training with DeepSpeed ZeRO-2 (Rajbhandari et al., 2020) and extend the open-r1 library (HuggingFace, 2025) to support PODS on multiple devices.
@@
We employ rule-based reward models that score rollouts, following standard practices in reasoning evaluation. Specifically, we reward an answer for correctness, format compliance, and the right number of thinking tags separately, resulting in a discrete but non-binary reward function. We provide additional details about the reward models Section˜A.1.
@@
In Section˜4.1, we compare the performance of GRPO and GRPO-PODS across six hardware and model settings listed in Table˜1. We show that for all the settings we test, GRPO-PODS consistently outperforms GRPO in terms of test performance throughout the training. Then, in Section˜4.2, we focus on setting (a), and analyze the effect of the rollout and update sizes $(n,m)$ on the performance of GRPO-PODS, providing empirical insights into how to choose the rollout and update sizes for GRPO-PODS. We also present additional experiments about different down-sampling rules (Section˜4.3) and different design choices of advantage normalization (Section˜A.3), additional evaluation results about PODS’ speed up ratio compared to GRPO (Section˜A.4) and test performance generalization across different test sets (Section˜A.5), and the average response length over the course of training (Section˜A.6).
(a) Training Qwen2.5 (3B) on GSM8K with one L40S GPU
(b) Training Llama3.2 (3B) on GSM8K with one L40S GPU
(c) Training Qwen2.5 (3B) on MATH with one L40S GPU
(d) Training Qwen2.5 (3B) on Chemistry with one L40S GPU
(e) Training Qwen2.5 (3B) on GSM8K with 8 H100 GPUs
(f) Training Qwen2.5 (7B) on GSM8K with 8 A100 GPUs
@@ 4.1 Comparing GRPO-PODS to baseline GRPO
We evaluate max-variance down-sampling PODS with GRPO against baseline GRPO using two experimental designs reflecting real-world constraints. For single-GPU settings (a-d), we compare against vanilla GRPO with the same training batch sizes $(m)$, where $m$ is selected so that one training batch fits within GPU memory. This corresponds to the comparison between the first and the third rows in Fig.˜2. For distributed settings (e-f), we compare against GRPO with gradient accumulation (GRPO-GA), the standard approach for scaling RLVR. In GRPO-GA, large training batches are processed through multiple gradient accumulation steps, enabling updating on larger effective batch sizes at the cost of increased communication overhead and iteration time. We fix the total rollouts generated per prompt $(n)$ and compare GRPO-GA against GRPO-PODS. This corresponds to the comparison between the second and the third rows in Fig.˜2. The detailed hyperparameters used for our experiments are listed in Section˜A.2.
Fig.˜3 shows test accuracy over wall-clock training time across all experiment configurations. PODS consistently achieves faster convergence: reaching the baselines’ peak accuracies at least $1.7\times$ faster (see Section˜A.4 for complete results) while often converging to a higher final performance. These results demonstrate PODS’ broad applicability across task domains (Math and Chemistry), model scales (3B-7B), architectures (Qwen2.5, Llama3.2), and deployment scenarios, making it a practical improvement for RLVR systems using GRPO.